# Optimal Control of a Subdiffusion model using Dirichlet-Neumann and Neumann-Neumann Waveform Relaxation Algorithms

### v1

@@ 5 Mesh Grading
We know that for fractional diffusion graded mesh in time give higher order accuracy than the uniform mesh, therefore one can consider as usual one sided graded mesh [9]. Now in control problem we need to solve both state and adjoint problem so the obvious extension of one sided graded mesh to both sided grading. Let $[0,T]$ be the total time domain and it is discretize into $2n:n\in\mathcal{N}$ non-overlapping sub-intervals by the nodes $t_{0}<t_{1}<\cdots<t_{2n}$. Now first we take usual graded mesh for time window $[0,T/2]$ using the nodes $t_{0}<t_{1}<\cdots<t_{n}$. Now for the remaining time window $[T/2,T]$ we take the similar grading but started from right end i.e. from $T$ using the nodes $t_{n+1}<t_{n+2}<\cdots<t_{2n}$. Hence for $j=0,1,\cdots,n$ time nodes $t_{j}=(j/n)^{(2-\alpha)/\alpha}T/2$ and $t_{2n-j}=T-t_{j}$.
6 Convergence of DNWR Algorithm
Now we analyze the convergence estimates of the DNWR algorithm (5)-(6), opting for the 1D model of the sub-diffusion and diffusion-wave equation to achieve better theoretical clarity and algebraic simplicity in our estimates. Our analysis focuses on a heterogeneous problem with $\kappa(\boldsymbol{x},t)=\kappa_{1}$ on $\Omega_{1}=(-h_{1},0)$ and $\kappa(\boldsymbol{x},t)=\kappa_{2}$ on $\Omega_{2}=(0,h_{2})$. We introduce the interface error $\omega^{(k)}(t)$ on state equation and $\upsilon^{(k)}(t)$ on adjoint equation and using the lemma 4.1 for transforming the Riemann-Liouville derivative to Caputo derivative the error euqation becomes:
@@ 6 Convergence of DNWR Algorithm
(19) $\upsilon^{(k)}(t)$ $=\phi p_{2}^{(k)}(0,t)+(1-\phi)\upsilon^{(k-1)}(t),$
Now we will analyze the system using semi-discrete analysis. We take the time nodes $0=t_{0}<t_{1}<\cdots<t_{n+1}=T$, $n\in\mathbb{N}$. Now we may take graded or uniform time mesh. Now if we use L1 scheme for left and right sided Caputo fractional derivative then for uniform and both sided graded mesh time discretize left and right fractional derivative matrix $A_{t}$ and $\bar{A}_{t}$ will be same for any $n$ in the case of uniform mesh and odd $n$ for both sided graded mesh. But for usual one sided graded mesh they will be different. This is because for uniform and both sided graded mesh follow the reflection symmetry i.e.$t_{j}=T-t_{n+1-j}$ for $j=0,1,\cdots,n+1$. This condition is violated for one sided graded mesh. As graded mesh itself take care of the fractional order so as the the fractional order $\alpha$ tends to $1$, the graded mesh also tends to the uniform mesh. With these mechanism of discretization we have
$A_{t}:=\begin{bmatrix}d_{1,1}&&&&\\ d_{2,2}-d_{2,1}&d_{2,1}&&&\\ \vdots&\ddots&\ddots&&\\ d_{n,n}-d_{n,n-1}&\cdots&&d_{n,1}&\\ d_{n+1,n+1}-d_{n+1,n}&\cdots&&d_{n+1,2}-d_{n+1,1}&d_{n+1,1}\end{bmatrix}.\textit{ And }B_{t}:=\begin{bmatrix}&&&1&\\ &&1&&\\ &\iddots&&\\ 1&&&&\\ &&&&\end{bmatrix}$
be the order $n+1$ square matrix coming from the source term in the right hand side. Here $d_{m,j}$ for $11\leq j\leq m\leq n+$ 1 are given in (13). For the sake of the analysis we choose only the uniform and both sided graded mesh and have the following state and adjoint equation:
(20) $\begin{array}[]{rcll}\begin{cases}A_{t}Y_{1}^{(k)}(\boldsymbol{x})=\kappa_{1}\partial_{xx}Y_{1}^{(k)}(\boldsymbol{x})-\frac{1}{\sigma}B_{t}P_{1}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(-h_{1},0),\\ Y_{1}^{(k)}(0)=\Omega^{(k-1)},\\ Y_{1}^{(k)}(-h_{1})=0,\end{cases}\end{array}\ \begin{array}[]{rcll}\begin{cases}A_{t}Y_{2}^{(k)}(\boldsymbol{x})=\kappa_{2}\partial_{xx}Y_{2}^{(k)}(\boldsymbol{x})-\frac{1}{\sigma}B_{t}P_{2}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(0,h_{2}),\\ \kappa_{2}\partial_{x}Y_{2}^{(k)}(0)=\kappa_{1}\partial_{x}Y_{1}^{(k)}(0),\\ Y_{2}^{(k)}(h_{2})=0,\end{cases}\end{array}$
@@
(21) $\begin{array}[]{rcll}\begin{cases}A_{t}P_{1}^{(k)}(\boldsymbol{x})=\kappa_{1}\partial_{xx}P_{1}^{(k)}(\boldsymbol{x})+B_{t}Y_{1}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(-h_{1},0),\\ P_{1}^{(k)}(0)=\Upsilon^{(k-1)},\\ P_{1}^{(k)}(-h_{1})=0,\end{cases}\end{array}\ \begin{array}[]{rcll}\begin{cases}A_{t}P_{2}^{(k)}(\boldsymbol{x})=\kappa_{2}\partial_{xx}P_{2}^{(k)}(\boldsymbol{x})+B_{t}Y_{2}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(0,h_{2}),\\ \kappa_{2}\partial_{x}P_{2}^{(k)}(0)=\kappa_{1}\partial_{x}P_{1}^{(k)}(0),\\ P_{2}^{(k)}(h_{2})=0,\end{cases}\end{array}$
@@
(23) $\Upsilon^{(k)}$ $=\phi P_{2}^{(k)}+(1-\phi)\Upsilon^{(k-1)},$
Here we have taken the discretization of $Y_{i}=[y_{i}^{1},y_{i}^{2},...,y_{i}^{n+1}]^{T}$ and $P_{i}$ into reverse direction i.e. $P_{i}=[p_{i}^{n},p_{i}^{n-1},...,p_{i}^{0}]^{T}$. For sake of analysis we combine both state and adjoint system into a single system and choosing the dependent variable $X_{i}=[Y_{i},P_{i}]^{T}$ Using the above two equations we form the system:
(24) $\begin{array}[]{rcll}\begin{cases}AX_{1}^{(k)}(\boldsymbol{x})=\kappa_{1}\partial_{xx}X_{1}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(-h_{1},0),\\ X_{1}^{(k)}(0)=\Pi^{(k-1)},\\ X_{1}^{(k)}(-h_{1})=0,\end{cases}\end{array}\ \begin{array}[]{rcll}\begin{cases}AX_{2}^{(k)}(\boldsymbol{x})=\kappa_{2}\partial_{xx}X_{2}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(0,h_{2}),\\ \kappa_{2}\partial_{x}X_{2}^{(k)}(0)=\kappa_{1}\partial_{x}X_{1}^{(k)}(0),\\ X_{2}^{(k)}(h_{2})=0,\end{cases}\end{array}$
Where, $A=\begin{bmatrix}A_{t}&\frac{1}{\sigma}B_{t}\\ -B_{t}&A_{t}\end{bmatrix}$. We choose $\Pi^{(k)}=[\Omega^{(k)},\Upsilon^{(k)}]^{T}$ for the combine form of the update condition for the new system and get:
@@
Proof 6.2.
We need to show that the numerical scheme is consistent and stable for the convergence of the solution. In discretization we have used L1-scheme on graded mesh or on uniform mesh, which is consistent. Now to show this scheme is stable we use the similar technique as in lemma 4.3. For that we first need the error equation of the semi discrete pde, which can be obtained by the substruction of the semi discrete pde with the continuous pde with the Taylor approximation of the time derivative. which is given as follows:
(26) $\begin{cases}AE_{1}(x)=\partial_{xx}E_{1}(x)+F(x),&\textrm{in}\;(-h_{1},0),\\ E_{1}(0)=0,\\ E_{1}(-h_{1})=0,\end{cases}$
@@
(27) $A\hat{E}_{1}=-\xi^{2}\hat{E}_{1}+\hat{F},$
$\hat{E}_{1}=(A+\xi^{2}I)^{-1}\hat{F}.$
Therefor the error depends on the absolute minimum eigenvalue of the matrix $(A+\xi^{2}I)$, which implies the eigenvalue of matrix $A$. Now $A=\begin{bmatrix}A_{t}&\frac{1}{\sigma}B_{t}\\ -B_{t}&A_{t}\end{bmatrix}$. From lemma 4.5 we know that eigenvalues of $A$ are the collection of eigenvalues of $A_{t}+\frac{i}{\sqrt{\sigma}}B_{t}$ and $A_{t}-\frac{i}{\sqrt{\sigma}}B_{t}$. Let, $A_{t}=\begin{bmatrix}A1_{t}&0\\ A2_{t}&a_{t}\end{bmatrix}$ be the order $n$ block matrix with $A1_{t}$ be the lower triangular block of order $(n-1)$ and $B_{t}=\begin{bmatrix}B1_{t}&0\\ 0&0\end{bmatrix}$ be the order $n$ block matrix with $B1_{t}$ be the anti-diagonal matrix of order $(n-1)$. From corollary 4.13 we know that each diagonal element from L1 scheme is positive, so $a_{t}\neq 0$. Again from the same corollary we have the matrix $A1_{t}$ is positive definite, hence from lemma 4.9 we conclude that the eigenvalues of $A_{t}+\frac{i}{\sqrt{\sigma}}B_{t}$ consist positive real part. In a similar manner we can also show the same for $A_{t}-\frac{i}{\sqrt{\sigma}}B_{t}$. Hence $(A+\xi^{2}I)$ will be a nonsingular matrix and the error will be bounded. Similar procedure can be used for the other part of (24) with Neumann boundary condition, where we have to use cosine function with certain phase difference.
Let, $A$ matrix is diagonalizable and $A=P\Lambda P^{-1}$. Solutions of (24) take the form:
@@
To ease the notional difficulties, define $a:=h_{1}/\sqrt{\kappa_{1}}\text{ and }b:=h_{2}/\sqrt{\kappa_{2}}$ for DNWR case. We first choose the most obvious case $a\neq b$, and define
@@
$+\begin{bmatrix}(1-(1+\sqrt{\kappa_{1}/\kappa_{2}})\theta)I&0\\ 0&(1-(1+\sqrt{\kappa_{1}/\kappa_{2}})\phi)I\end{bmatrix}\Pi^{(k-1)}.$
@@
Theorem 6.3.
If for every eigenvalue $\lambda_{i}$ of $A$, the DNWR algorithm for optimal control converge if the quantity
$\left|\frac{\sinh((b-a)\sqrt{\lambda_{i}})}{\sinh(a\sqrt{\lambda_{i}})\cosh(b\sqrt{\lambda_{i}})}\right|<\frac{\sqrt{\kappa_{1}}+\sqrt{\kappa_{2}}}{\sqrt{\kappa_{1}}}.$
Lemma 6.4.
For any $\alpha,\beta$ nonzero positive real number and $z\in\mathbb{C}$
$\left|\frac{\exp(-\alpha z)}{1-\exp(-\beta z)}\right|\leq\frac{\exp(-\alpha\Re(z))}{1-\exp(-\beta\Re(z))}.$
Proof 6.5.
Let $z=x+iy$. Then
$\left|\frac{\exp(-\alpha z)}{1-\exp(-\beta z)}\right|$ $=\frac{\exp(-\alpha x)}{\sqrt{(1-\exp(-\beta x)\cos(\beta y))^{2}+(\exp(-\beta x)\sin(\beta y))^{2}}}$
$\leq\frac{\exp(-\alpha x)}{\sqrt{1-2\exp(-\beta x)\cos(\beta y)+\exp(-2\beta x)}}$
@@
Theorem 6.6.
Let $\lambda_{i}$ are the eigenvalues and $P$ is the eigenvector matrix of matrix $A$. If $\lambda=\min(\Re(\sqrt{\lambda_{i}}))$, then the interface error term of the DNWR algorithm satisfies the following bound:
@@
Proof 6.7.
Left multiplying with $P^{-1}$ in (31) error equation becomes:
@@
$=-\frac{\sqrt{\kappa_{1}}}{\sqrt{\kappa_{1}}+\sqrt{\kappa_{2}}}\frac{\sinh((b-a)\sqrt{\Lambda})}{\sinh(b\sqrt{\Lambda})\cosh(a\sqrt{\Lambda})}P^{-1}\Pi^{(k-1)}$
Now consider the $i^{th}$ entry in (32) we have:
(33) $\left(P^{-1}\Pi^{(k)}\right)_{i}$ $=-\frac{\sqrt{\kappa_{1}}}{\sqrt{\kappa_{1}}+\sqrt{\kappa_{2}}}\frac{\sinh((b-a)\sqrt{\lambda_{i}})}{\sinh(b\sqrt{\lambda_{i}})\cosh(a\sqrt{\lambda_{i}})}\left(P^{-1}\Pi^{(k-1)}\right)_{i}$
Let $(W)_{ii}:=\frac{\sinh((b-a)\sqrt{\lambda_{i}})}{\sinh(b\sqrt{\lambda_{i}})\cosh(a\sqrt{\lambda_{i}})}$ be the weight. And we will estimate the weight by simplifying the hyperbolic terms in terms of exponential and we have:
@@
(35) $|I_{1}|$ $=2\left|\frac{\exp(-a\sqrt{\lambda_{i}})}{\left(1-\exp(-2a\sqrt{\lambda_{i}})\right)}\right|\left|\frac{\exp(-a\sqrt{\lambda_{i}})}{\left(1+\exp(-2b\sqrt{\lambda_{i}})\right)}\right|$
Using the inequality from lemma 6.4, we have
$|I_{1}|$ $\leq 2\frac{\exp(-2a\Re(\sqrt{\lambda_{i}}))}{\left(1-\exp(-2a\Re(\sqrt{\lambda_{i}}))\right)\left(1-\exp(-2b\Re(\sqrt{\lambda_{i}}))\right)}$
Further using $\lambda=\min(\Re(\sqrt{\lambda_{i}}))$, we obtain
$|I_{1}|$ $\leq 2\frac{\exp(-2a\lambda))}{\left(1-\exp(-2a\lambda))\right)\left(1-\exp(-2b\lambda))\right)}$
By similar argument
@@
$\leq 2\frac{\exp(-2b\lambda)}{\left(1-\exp(-2a\lambda)\right)\left(1-\exp(-2b\lambda)\right)}$
@@
$\leq\frac{\cosh((b-a)\lambda)}{\sinh(a\lambda)\sinh(b\lambda)}$
Hence $\|W\|_{\infty}\leq\frac{\cosh((b-a)\lambda)}{\sinh(a\lambda)\sinh(b\lambda)}$. Now taking the supremum norm on both side in (37) we have:
@@
$\leq\left(\frac{\sqrt{\kappa_{1}}}{\sqrt{\kappa_{1}}+\sqrt{\kappa_{2}}}\right)^{k}\left(\frac{\cosh((b-a)\lambda)}{\sinh(a\lambda)\sinh(b\lambda)}\right)^{k}\left\|P^{-1}\Pi^{(0)}\right\|_{\infty}$
Again using the norm inequality we have the required result:
@@
7 Convergence of NNWR Algorithm
To study the convergence of the NNWR algorithm (LABEL:NNWRD)-(11) in 1D, we first divide the spatial interval $(-L_{1},L_{2})$ into $N$ subintervals $(x_{i-1},x_{i})$ with no overlap or cross point, where $i=1,2,...,N$. We define $a_{i}$ as the length of each subinterval, and we assume different constant diffusion coefficients $\kappa_{i}$ in each subinterval $\Omega_{i}$. We also define $\omega_{i}^{(k)}(t)$ as the initial error along the artificial boundary at ${x_{i}}_{i=1}^{N-1}$ and set $\omega_{0}^{(k)}(t)=0$ at $x=x_{0}$ and $\omega_{N}^{(k)}(t)=0$ at $x=x_{N}$ for the original boundaries. By applying the Laplace transform in time, we can simplify the equations (LABEL:NNWRD)-(11) to

### v2

@@ 6 Mesh Grading
We are aware that utilizing a fractional diffusion graded mesh over time yields a higher order of accuracy, denoted as $\mathcal{O}(M^{-(2-\beta)})$, compared to a uniform mesh, which offers an accuracy of $\mathcal{O}(M^{-\beta})$, where $\beta$ represents fractional order and $M$ denotes number of nodes. Thus, it is customary to adopt a one-sided graded mesh, as suggested in [stynes2017error].
In control problems, solving both the state and adjoint problems necessitates extending the one-sided graded mesh approach to a both-sided grading. Let’s consider a total time domain $[0,T]$, divided into $n+1$ non-overlapping sub-intervals, with $n$ being odd, defined by the nodes $0=t_{0}<t_{1}<\cdots<t_{n+1}=T$. Initially, we employ the usual graded mesh for the time window $[0,T/2]$, utilizing nodes $t_{0}<t_{1}<\cdots<t_{(n+1)/2}$. Subsequently, for the remaining time window $[T/2,T]$, we apply a similar grading, starting from the right end, i.e., from $T$, using nodes $t_{n+1}>t_{n}>\cdots>t_{(n+1)/2}$. Thus, for $j=0,1,\ldots,(n+1)/2$, the time nodes are defined as $t_{j}=(j/(n+1)/2)^{(2-\alpha)/\alpha}T/2$ and $t_{n+1-j}=T-t_{j}$.
7 Convergence of Dirichlet-Neumann Waveform Relaxation
Presently, we analyze the convergence assessments of the DNWR algorithm (5)-(6), selecting the 1D model of the sub-diffusion and diffusion equation for enhanced theoretical simplicity and algebraic straightforwardness in our evaluations. Our analysis is directed towards a heterogeneous scenario where $\kappa(\boldsymbol{x},t)=\kappa_{1}$ over $\Omega_{1}=(-h_{1},0)$ and $\kappa(\boldsymbol{x},t)=\kappa_{2}$ over $\Omega_{2}=(0,h_{2})$. We introduce the interface error $\omega^{(k)}(t)$ on state equation and $\upsilon^{(k)}(t)$ on adjoint equation and using the lemma 1 for transforming the Riemann-Liouville derivative to Caputo derivative in the error equation to obtain:
@@ 7 Convergence of Dirichlet-Neumann Waveform Relaxation
(20) $\upsilon^{(k)}(t)$ $=\phi p_{2}^{(k)}(0,t)+(1-\phi)\upsilon^{(k-1)}(t).$
Now, we will conduct an analysis of the system using semi-discrete methods. We define time nodes as $0=t_{0}<t_{1}<\cdots<t_{n+1}=T$, where $n$ belongs to the set of natural numbers. Here, we have the option to employ either a graded or uniform time mesh. When employing the L1 scheme for both left and right-sided Caputo fractional derivatives, the discretized matrices for these derivatives will be the same in uniform mesh for any $n\in\mathbb{N}$ and in both sides of a graded mesh when $n$ is odd. However, for the usual one-sided graded mesh, they will differ. This is because both the uniform and two-sided graded meshes adhere to a reflective symmetry, i.e., $t_{j}=T-t_{n+1-j}$ for $j=0,1,\cdots,n+1$. This symmetry condition is not satisfied for the one-sided graded mesh. As the graded mesh inherently accounts for the fractional order, as the fractional order $\alpha$ tends towards $1$, the graded mesh converges towards the uniform mesh. With these discretization mechanisms in place, we can proceed with the analysis. Let, the discretized time fractional derivative defined as
$L:=\begin{bmatrix}d_{1,1}&&&&\\ d_{2,2}-d_{2,1}&d_{2,1}&&&\\ \vdots&\ddots&\ddots&&\\ d_{n,n}-d_{n,n-1}&\cdots&&d_{n,1}&\\ d_{n+1,n+1}-d_{n+1,n}&\cdots&&d_{n+1,2}-d_{n+1,1}&d_{n+1,1}\end{bmatrix}.\textit{ And }\bar{J}:=\begin{bmatrix}&&&1&\\ &&1&&\\ &\iddots&&\\ 1&&&&\\ &&&&\end{bmatrix}$
be the order $n+1$ square matrix coming from the source term in the right hand side. Here $d_{m,j}$ for $1\leq j\leq m\leq n+1$ are given in (13). For the sake of the analysis we choose only the uniform and both sided graded mesh and have the following state and adjoint equation:
(21) $\begin{array}[]{rcll}\begin{cases}LY_{1}^{(k)}(\boldsymbol{x})=\kappa_{1}\partial_{xx}Y_{1}^{(k)}(\boldsymbol{x})-\frac{1}{\sigma}\bar{J}P_{1}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(-h_{1},0),\\ Y_{1}^{(k)}(0)=\Omega^{(k-1)},\\ Y_{1}^{(k)}(-h_{1})=0,\end{cases}\end{array}\ \begin{array}[]{rcll}\begin{cases}LY_{2}^{(k)}(\boldsymbol{x})=\kappa_{2}\partial_{xx}Y_{2}^{(k)}(\boldsymbol{x})-\frac{1}{\sigma}\bar{J}P_{2}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(0,h_{2}),\\ \kappa_{2}\partial_{x}Y_{2}^{(k)}(0)=\kappa_{1}\partial_{x}Y_{1}^{(k)}(0),\\ Y_{2}^{(k)}(h_{2})=0,\end{cases}\end{array}$
@@
(22) $\begin{array}[]{rcll}\begin{cases}LP_{1}^{(k)}(\boldsymbol{x})=\kappa_{1}\partial_{xx}P_{1}^{(k)}(\boldsymbol{x})+\bar{J}Y_{1}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(-h_{1},0),\\ P_{1}^{(k)}(0)=\Upsilon^{(k-1)},\\ P_{1}^{(k)}(-h_{1})=0,\end{cases}\end{array}\ \begin{array}[]{rcll}\begin{cases}LP_{2}^{(k)}(\boldsymbol{x})=\kappa_{2}\partial_{xx}P_{2}^{(k)}(\boldsymbol{x})+\bar{J}Y_{2}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(0,h_{2}),\\ \kappa_{2}\partial_{x}P_{2}^{(k)}(0)=\kappa_{1}\partial_{x}P_{1}^{(k)}(0),\\ P_{2}^{(k)}(h_{2})=0,\end{cases}\end{array}$
@@
(24) $\Upsilon^{(k)}$ $=\phi P_{2}^{(k)}+(1-\phi)\Upsilon^{(k-1)}.$
Here we have taken the discretization of $Y_{i}:=[y_{i}^{1},y_{i}^{2},...,y_{i}^{n+1}]^{T}$ in forward direction and $P_{i}$ into reverse direction i.e. $P_{i}:=[p_{i}^{n},p_{i}^{n-1},...,p_{i}^{0}]^{T}$. For sake of analysis we combine both state and adjoint system into a single system and choosing the dependent variable $X_{i}:=[Y_{i};P_{i}]$ Using the above two equations we form the system:
(25) $\begin{array}[]{rcll}\begin{cases}\mathbb{L}X_{1}^{(k)}(\boldsymbol{x})=\kappa_{1}\partial_{xx}X_{1}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(-h_{1},0),\\ X_{1}^{(k)}(0)=\Pi^{(k-1)},\\ X_{1}^{(k)}(-h_{1})=0,\end{cases}\end{array}\ \begin{array}[]{rcll}\begin{cases}\mathbb{L}X_{2}^{(k)}(\boldsymbol{x})=\kappa_{2}\partial_{xx}X_{2}^{(k)}(\boldsymbol{x}),&\textrm{in}\;(0,h_{2}),\\ \kappa_{2}\partial_{x}X_{2}^{(k)}(0)=\kappa_{1}\partial_{x}X_{1}^{(k)}(0),\\ X_{2}^{(k)}(h_{2})=0,\end{cases}\end{array}$
where, $\mathbb{L}:=\begin{bmatrix}L&\frac{1}{\sigma}\bar{J}\\ -\bar{J}&L\end{bmatrix}$. We choose $\Pi^{(k)}:=[\Omega^{(k)};\Upsilon^{(k)}]$ for the combine form of the update condition for the new system and get:
@@
Proof.
We need to show that the numerical scheme is consistent and stable for the convergence of the solution. In discretization we have used L1-scheme on graded mesh or on uniform mesh, which is consistent. Now to show this scheme is stable we use the similar technique as in lemma 2. For that we first need the error equation of the semi discrete pde, which can be obtained by the substruction of the semi discrete pde with the continuous pde with the Taylor approximation of the time derivative. which is given as follows:
(27) $\begin{cases}\mathbb{L}E_{1}(x)=\partial_{xx}E_{1}(x)+F(x),&\textrm{in}\;(-h_{1},0),\\ E_{1}(0)=0,\\ E_{1}(-h_{1})=0,\end{cases}$
@@
(28) $\mathbb{L}\hat{E}_{1}=-\xi^{2}\hat{E}_{1}+\hat{F}$,
$\hat{E}_{1}=(\mathbb{L}+\xi^{2}\mathbb{I})^{-1}\hat{F}$.
Therefor the error is contingent upon the smallest absolute eigenvalue of the matrix $(\mathbb{L}+\xi^{2}\mathbb{I})$. Now $\mathbb{L}=\begin{bmatrix}L&\frac{1}{\sigma}\bar{J}\\ -\bar{J}&L\end{bmatrix}$. From lemma 3 we know that eigenvalues of $\mathbb{L}$ are the collection of eigenvalues of $L+\frac{i}{\sqrt{\sigma}}\bar{J}$ and $L-\frac{i}{\sqrt{\sigma}}\bar{J}$. Let, $L=\begin{bmatrix}L_{1}&0\\ L_{2}&d_{n+1,1}\end{bmatrix}$ be the order $(n+1)$ block matrix with $L_{1}$ be the lower triangular block of order $n$ and $\bar{J}=\begin{bmatrix}J&O\\ O&0\end{bmatrix}$ be the order $(n+1)$ block matrix with $J$ be the exchange matrix of order $n$. From corollary 6 we know that each diagonal element from L1 scheme is positive, so $d_{n+1,1}>0$. Again from the same corollary we have the matrix $L_{1}$ is positive definite, hence from lemma 4 we conclude that the eigenvalues of $L+\frac{i}{\sqrt{\sigma}}\bar{J}$ consist positive real part. In a similar manner we can also show the same for $L-\frac{i}{\sqrt{\sigma}}\bar{J}$. Hence $(\mathbb{L}+\xi^{2}\mathbb{I})$ be a nonsingular matrix and the error is bounded. Similar procedure can be used for the other part of (25) with Neumann boundary condition, where we have to use cosine function with certain phase difference. ∎
Let the matrix $\mathbb{L}$ be diagonalizable by the eigenvector matrix $P$ corresponding to the eigenvalue matrix $\Lambda$, which means $\mathbb{L}=P\Lambda P^{-1}$. For the sake of notation simplicity, $f(\Lambda)$, where $f$ is any scalar-valued function, implies that the function $f$ operates solely on the diagonal elements of $\Lambda$ in a pointwise manner, while all other elements are zero. Solutions of (25) take the form:
@@
For notational simplicity, consider $a:=h_{1}/\sqrt{\kappa_{1}},b:=h_{2}/\sqrt{\kappa_{2}}$. Then for $a\neq b$ (30) can be written as
@@
$\hskip 56.9055pt+\begin{bmatrix}(1-(1+\sqrt{\kappa_{1}/\kappa_{2}})\theta)I&0\\ 0&(1-(1+\sqrt{\kappa_{1}/\kappa_{2}})\phi)I\end{bmatrix}\Pi^{(k-1)}.$
@@
Lemma 12.
For any $\alpha,\beta$ nonzero positive real number and $z\in\mathbb{C}$ with $\Re(z))>0$
$\left|\frac{\exp(-\alpha z)}{1\pm\exp(-\beta z)}\right|\leq\frac{\exp(-\alpha\Re(z))}{1-\exp(-\beta\Re(z))}.$
Proof.
Let $z=x+iy,x>0$. Then
$\left|\frac{\exp(-\alpha z)}{1\pm\exp(-\beta z)}\right|$ $=\frac{\exp(-\alpha x)}{\sqrt{(1\pm\exp(-\beta x)\cos(\beta y))^{2}+(\exp(-\beta x)\sin(\beta y))^{2}}}$
$=\frac{\exp(-\alpha x)}{\sqrt{1\pm 2\exp(-\beta x)\cos(\beta y)+\exp(-2\beta x)}}$
@@
Theorem 13.
Let time discretize combined state and adjoint matrix $\mathbb{L}$ in (25) be diagonalizable by the eigenvector matrix $P$ corresponding to the eigenvalues $\{\lambda_{i}:i=1,\ldots,2n+2\}$. If $\lambda=\min(\Re(\sqrt{\lambda_{i}}))$, then the interface error term of the DNWR algorithm in (5)-(6) in 1D for relaxation parameters $\theta=\phi=1/(1+\sqrt{\kappa_{1}/\kappa_{2}})$ satisfies the following bound
@@
Proof.
Let $\mathbb{L}$ be diagonalizable by eigenvector matrix $P$ corresponding to the eigenvalue matrix $\Lambda$, i.e. $\mathbb{L}=P\Lambda P^{-1}$. We know from the proof of Theorem 11 that all the eigenvalues of $\mathbb{L}$ lies in first or fourth quadrant in complex plane therefore the principal square roots of eigenvalues also lie in first or fourth quadrant, hence has positive real part i.e. $\lambda=\min(\Re(\sqrt{\lambda_{i}}))$ is well defined.
Now for the proof we take the update part (32) then left multiplying on both side with $P^{-1}$ to have
@@
$=-\frac{\sqrt{\kappa_{1}}}{\sqrt{\kappa_{1}}+\sqrt{\kappa_{2}}}\frac{\sinh((b-a)\sqrt{\Lambda})}{\sinh(b\sqrt{\Lambda})\cosh(a\sqrt{\Lambda})}P^{-1}\Pi^{(k-1)}=:-\frac{\sqrt{\kappa_{1}}}{\sqrt{\kappa_{1}}+\sqrt{\kappa_{2}}}WP^{-1}\Pi^{(k-1)}.$
Let $(W)_{ii}:=\frac{\sinh((b-a)\sqrt{\lambda_{i}})}{\sinh(b\sqrt{\lambda_{i}})\cosh(a\sqrt{\lambda_{i}})}$ be the weight. And we will estimate the weight by simplifying the hyperbolic terms in terms of exponential, which is as follows
@@
(35) $|I_{1}|$ $=2\left|\frac{\exp(-a\sqrt{\lambda_{i}})}{\left(1-\exp(-2a\sqrt{\lambda_{i}})\right)}\right|\left|\frac{\exp(-a\sqrt{\lambda_{i}})}{\left(1+\exp(-2b\sqrt{\lambda_{i}})\right)}\right|.$
Using the inequality from lemma 12, we have
$|I_{1}|$ $\leq 2\frac{\exp(-2a\Re(\sqrt{\lambda_{i}}))}{\left(1-\exp(-2a\Re(\sqrt{\lambda_{i}}))\right)\left(1-\exp(-2b\Re(\sqrt{\lambda_{i}}))\right)}.$
Further using $\lambda=\min(\Re(\sqrt{\lambda_{i}}))$ and the property of monotonic decreasing function, we obtain
$|I_{1}|$ $\leq 2\frac{\exp(-2a\lambda))}{\left(1-\exp(-2a\lambda))\right)\left(1-\exp(-2b\lambda))\right)}.$
By similar argument we have the bound
@@
$\leq 2\frac{\exp(-2b\lambda)}{\left(1-\exp(-2a\lambda)\right)\left(1-\exp(-2b\lambda)\right)}.$
@@
$\leq\frac{\cosh((b-a)\lambda)}{\sinh(a\lambda)\sinh(b\lambda)}.$
Hence $\|W\|_{\infty}\leq\frac{\cosh((b-a)\lambda)}{\sinh(a\lambda)\sinh(b\lambda)}$. Now taking the supremum norm on both side in (33) we have
@@
$\leq\left(\frac{\sqrt{\kappa_{1}}}{\sqrt{\kappa_{1}}+\sqrt{\kappa_{2}}}\right)^{k}\left(\frac{\cosh((b-a)\lambda)}{\sinh(a\lambda)\sinh(b\lambda)}\right)^{k}\left\|P^{-1}\Pi^{(0)}\right\|_{\infty}.$
Again using the norm inequality we have the required result
@@
8 Convergence of Neumann-Neumann Waveform Relaxation
To analyze the convergence for the system described by equations (9)-(10) in 1D setting, we initially partition the domain $(L_{1},L_{2})$ into $N$ non-overlapping subdomains $(x_{i-1},x_{i})$, where i=1,2,,Ni=1,2,...,N. Each subdomain is characterized by a length $a_{i}$, and we consider distinct constant generalized diffusion coefficients $\kappa_{i}$ within each subdomain $\Omega_{i}$. Additionally, we introduce $\omega_{i}^{(k)}(t)$ and $\upsilon_{i}^{(k)}(t)$ to represents the error terms at ${x_{i}}_{i=1}^{N-1}$ for the state and adjoint equations, respectively. Than set boundary conditions as $\omega_{0}^{(k)}(t)=0,\upsilon_{0}^{(k)}(t)=0$ at $x=x_{0}$, and $\omega_{N}^{(k)}(t)=0,\upsilon_{N}^{(k)}(t)=0$ at $x=x_{N}$. Now using the Lemma 1 to convert the Riemann-Liouville derivative into the Caputo derivative within the error equation yields: